# Spectral Deconvolution without the Deconvolution: Extracting Temperature from X-ray Thomson Scattering Spectra without the Source-and-Instrument Function

### v1

@@ III.3 Robustness to Spectral Noise
Of course, the results will also depend on the number of photons collected in a given spectrum. The spectra shown in Fig. 3 contain $\sim 3\times 10^{6}$ photons in total. At XFEL facilities, the high repetition rate of the XFEL (up to 10 Hz per frame at the European XFEL, containing potentially multiple bunches at 4.5 MHz [68]; and up to 120 Hz per frame at the LCLS [71]), means that it is possible to collect tens of thousands of shots very quickly. The high degree of stability of XFEL pulses, combined with recent developments in stable high repetition rate (1-10 Hz) laser drive systems such as DiPOLE-100X [72] and ReLaX [73], means that collecting the sorts of photon statistics shown here is achievable. In the case of backlighter sources, such as those used at the National Ignition Facility [4, 56] and OMEGA laser facility [2, 43], the laser heating of the foil results in a tremendous amount of photon emission, such that good photon statistics can be accumulated in single shots (a necessity, given a delay of several hours between shots).
@@
For $T=20$ eV, when the number of photons is reduced by a factor of 10, temperatures may still be inferred from the different ratios, however there are now differences in the three inferred temperatures, and the uncertainty on these temperatures is increased to $\sim 2$ eV. Within the uncertainty indicated by the spread of the convergence regions, there is overlap between the temperatures. However, the results also likely have some dependence on the specific spectrum sampled due to the increased photon counting noise. It would also be challenging to infer any non-equilibrium effects unless there are temperature discrepancies of at least 10 eV. Reducing the photon counts by another factor of 10 reduces the upshifted signal sufficiently that detailed balance cannot be probed, and a temperature cannot be extracted. There is a completely flat region at $x>190$ eV for the 201020/10 ratio, however this convergence is spurious, and due to the low photon detections at large energy shifts. Additionally, an accurate, converged inferred temperature would not be expected to suddenly emerge after the collapse of the ratio seen between x=120x=120–190 eV. Therefore this apparent temperature can be ignored.
@@
Once again, the results shown in this section strongly depend on the Rayleigh weight model: reducing the Rayleigh weight to increase the portion of photons scattered inelastically means that fewer total photons can suffice for a given temperature, and the temperature discrepancies required to infer non-equilibrium are likewise lowered. Since the ratio of inelastic to elastic scattering depends on the specific system being probed, planning is required to determine the number of photons that must be detected and the optimal choice of scattering angles in order to adequately infer the temperature or non-equilibrium from the ITCF ratios.
@@ III.5.1 Mosaicity
Figure 14: Temperature convergences for $T=20$ eV for three different scattering angle ratios, but with the mosaicity of one of the crystals different from the original 0.063: 0.07 (red), 0.08 (blue), 0.09 (orange), and 0.10 (green). Also shown in black is the original case where the crystals are identical (i.e. 0.063/0.063). The ratios corresponding to each curve in the legend indicate which mosaicity is in the numerator and denominator.
@@
In Fig. 14 and 15, we compare inferred temperature convergence for $T=20$ eV and $T=50$ eV, respectively, where one of the spectrometers has the original 0.063 mosaicity, and the other has increasing mosaicity. Both the case where the denominator spectrum has the increasing mosaicity and the case where the numerator has the increasing mosaicity are plotted. This distinction has a noticeable effect on the convergence behaviour for $T=20$ eV ratios and for the 201020/10 ratio at $T=50$ eV – these are the cases where the elastic scattering is very strong compared to the inelastic scattering. Once again, when the elastic scattering is very prominent, it has a noticeable effect on the convergence due to differences in the 8.5 keV SIF between the two ratios. Removing the elastic feature while keeping the inelastically scattered signal the same results in excellent agreement between the 15010150/10 and 15020150/20 ratios regardless of the mosaicity and their position in the ratio, as shown in Fig. 16; a similar observation is made for the $T=50$ eV case in Fig. 15. The 201020/10 case in Fig. 15 and 16 shows generally poor convergence stability to changes in the mosaicity, indicating that the narrow plasmon peak shapes are strongly influenced by the shape of the SIFs in their vicinity.
@@ III.5.2 Intrinsic Rocking Curve Widths
Figure 17: Temperature convergences for $T=20$ eV for three different scattering angle ratios, but with the IRC width of one of the crystals changed from the original by a factor of 1.05 (red), 1.1 (blue), 1.5 (orange), and 2.0 (green). Also shown in black is the original case where both crystals are identical. The ratios corresponding to each curve in the legend indicate which IRC width is in the numerator and denominator.
@@
The results are plotted in Fig. 17 and 18 for $T=20$ eV and 50 eV, respectively. We consider both the case that the broader IRC is in the numerator of the ratio and in the denominator. As with the mosaicity, the situations in which the elastic peak is most prominent (i.e. low temperature and low scattering angles), there is a fairly strong sensitivity to the width of the IRC, however the sensitivity is much lower than for the mosaicity. The $T=50$ eV case shows consistent convergence across all sets of ratios considered. Indeed, even using the standard Rayleigh weight model at $T=20$ eV, the 15010150/10 and 15020150/20 ratios give fairly consistent temperatures, especially when the broader IRC crystal is used to measure the low scattering angle spectrum. This points to the fact that planning can be done to decide which crystals of a set should be used to measure the low angle spectrum and the high angle spectrum for optimally using the ratio method to extract temperature.
@@ Acknowledgments
This work was partially supported by the Center for Advanced Systems Understanding (CASUS), financed by Germany’s Federal Ministry of Education and Research (BMBF) and the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Union’s Just Transition Fund (JTF) within the project Röntgenlaser-Optimierung der Laserfusion (ROLF), contract number 5086999001, co-financed by the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2022 research and innovation programme (Grant agreement No. 101076233, “PREXTREME”). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. A.K. and A.D.B. were partially supported by the U.S. Department of Energy Science Campaign 1. Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is co-authored by employees of NTESS. The employees, not NTESS, own the right, title and interest in and to the written work and are responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan. The work of M.P.B. was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344 and supported by Laboratory Directed Research and Development (LDRD) Grant No. 25-ERD-047. The authors gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy/Grete at NHR-Nord@Göttingen as part of the NHR infrastructure. The calculations for this research were conducted with computing resources under the project mvp00024.

### v2

@@ III.3 Robustness to Spectral Noise
The results will also depend on the number of photons collected in a given spectrum. The spectra shown in Fig. 3 contain $\sim 3\times 10^{6}$ photons in total. At XFEL facilities, the high repetition rate of the XFEL (up to 10 Hz per frame at the European XFEL, containing potentially multiple bunches at 4.5 MHz [81]; and up to 120 Hz per frame at the LCLS [40]), means that it is possible to collect tens of thousands of shots very quickly. The high degree of stability of XFEL pulses, combined with recent developments in stable high repetition rate (1-10 Hz) laser drive systems such as DiPOLE-100X [62] and ReLaX [59], means that collecting the sorts of photon statistics shown here is achievable. In the case of backlighter sources, such as those used at the National Ignition Facility [16, 17] and OMEGA laser facility [28, 56], the laser heating of the foil results in a tremendous amount of photon emission, such that good photon statistics can be accumulated in single shots (a necessity, given a delay of several hours between shots).
@@
For $T=20$ eV, when the number of photons is reduced by a factor of 10, temperatures may still be inferred from the different ratios, however there are now differences in the three inferred temperatures, and the uncertainty on these temperatures is increased to $\sim 2$ eV. Within the uncertainty indicated by the spread of the convergence regions, there is overlap between the temperatures. However, the results also likely have some dependence on the specific spectrum sampled due to the increased photon counting noise. It would also be challenging to infer any non-equilibrium effects unless there are temperature discrepancies of at least 10 eV. Reducing the photon counts by another factor of 10 reduces the upshifted signal sufficiently that detailed balance cannot be probed, and a temperature cannot be extracted. There is a completely flat region at $x>190$ eV for the 201020/10 ratio, however this convergence is spurious and is because of the low photon detections at large energy shifts. Additionally, an accurate, converged inferred temperature would not be expected to suddenly emerge after the collapse of the ratio seen between x=120x=120–190 eV. Therefore this apparent temperature can be ignored.
@@
Once again, the results shown in this section strongly depend on the Rayleigh weight: reducing the Rayleigh weight to increase the portion of photons scattered inelastically means that fewer total photons can suffice for a given temperature, and the temperature discrepancies required to infer non-equilibrium are likewise lowered. Since the ratio of inelastic to elastic scattering depends on the specific system being probed, planning is required to determine the number of photons that must be detected and the optimal choice of scattering angles in order to adequately infer the temperature or non-equilibrium from the ITCF ratios.
@@ III.5.1 Mosaicity
Figure 15: Temperature convergences for $T=20$ eV for three different scattering angle ratios, but with the mosaicity of one of the crystals different from the original 0.063: 0.07 (red), 0.08 (blue), 0.09 (orange), and 0.10 (green). Also shown in black is the original case where the crystals are identical (i.e. 0.063/0.063). The top plots show the cases when the higher mosaicity is in the denominator, and the bottom plots when it is in the numerator.
@@
In Fig. 15 and 16, we compare inferred temperature convergence for $T=20$ eV and $T=50$ eV, respectively, where one of the spectrometers has the original 0.063 mosaicity, and the other has increasing mosaicity. Both the case where the denominator spectrum has the increasing mosaicity and the case where the numerator has the increasing mosaicity are plotted. This distinction has a noticeable effect on the convergence behaviour for $T=20$ eV ratios and for the 201020/10 ratio at $T=50$ eV – these are the cases where the elastic scattering is very strong compared to the inelastic scattering. Once again, when the elastic scattering is very prominent, it has a noticeable effect on the convergence due to differences in the 8.5 keV SIF between the two ratios. Removing the elastic feature while keeping the inelastically scattered signal the same results in excellent agreement between the 15010150/10 and 15020150/20 ratios regardless of the mosaicity and their position in the ratio, as shown in Fig. 17; a similar observation is made for the $T=50$ eV case in Fig. 16. The 201020/10 case in Fig. 16 and 17 shows generally poor convergence stability to changes in the mosaicity, indicating that the narrow plasmon peak shapes are strongly influenced by the shape of the SIFs in their vicinity.
@@ III.5.2 Intrinsic Rocking Curve Widths
Figure 18: Temperature convergences for $T=20$ eV for three different scattering angle ratios, but with the IRC width of one of the crystals changed from the original by a factor of 1.05 (red), 1.1 (blue), 1.5 (orange), and 2.0 (green). Also shown in black is the original case where both crystals are identical. The top plots show the cases when the wider IRC is in the denominator, and the bottom plots when it is in the numerator.
@@
The results are plotted in Fig. 18 and 19 for $T=20$ eV and 50 eV, respectively. We consider both the case that the broader IRC is in the numerator of the ratio and in the denominator. As with the mosaicity, the situations in which the elastic peak is most prominent (i.e. low temperature and low scattering angles), there is a fairly strong sensitivity to the width of the IRC, however the sensitivity is much lower than for the mosaicity. The $T=50$ eV case shows consistent convergence across all sets of ratios considered. Indeed, even using the standard Rayleigh weight model at $T=20$ eV, the 15010150/10 and 15020150/20 ratios give fairly consistent temperatures, especially when the broader IRC crystal is used to measure the low scattering angle spectrum. This points to the fact that planning can be done to decide which crystals of a set should be used to measure the low angle spectrum and the high angle spectrum for optimally using the ratio method to extract temperature.
@@
III.6 Application to Laser Facility Experiments
Figure 21: (a) Ray traced instrument functions for 8, 8.5, 9, 9.5, and 10 keV photons on the MACS spectrometer, showing the variation in the instrument function (especially the peak reflectivity) across the spectrometer. (b) Ray traced XRTS spectra of compressed Be at conditions reported in Ref. [16] at three different scattering angles. Additionally, the elastic scattering feature is shown to indicate the form of the SIF broadening. (c) Temperature convergences for the three different ratios.
[…]
This setup presents a number of potential challenges for the method. First, the SIF broadening is substantially larger, which will require a much larger integration range to remove the SIF broadening than before. However, the MACS spectrometer has a very large spectral range of >3>3 keV, which will provide more than sufficient integration range to observe the convergence of the inferred temperature.
Second, and more problematic, is that the SIF varies substantially across the spectrometer, as shown in Fig. 21 (a). This is due to the out-of-focus geometry of the MACS spectrometer, which results in photons spread across conical arcs centred around the focused energy of 8.6 keV. Evidently this raises questions as to the applicability of the convolution approximation in Eqs. (3) and (8). Nevertheless, this approximation routinely used in analysis [16, 22], and we shall still assume it to hold here.
Third is that the large capsule size, source profile, and finite-size of the spectrometer would mean XRTS spectra over a range of scattering vectors would be measured on the detector. Moreover, conditions within the capsule are not uniform, with substantial variations in temperature and density predicted by radiative hydrodynamics modelling [16]. However, such considerations fall outside the scope of the current work, so we treat conditions within the capsule to be uniform. Still, these effects should be examined in detail in future works.
For the DSF, we model laser-compressed Be using the conditions reported in Row 1 of Table 1 in Ref. [16]; a mass density of 6.7 g cm-3, ionization state of 3, and a temperature of 110 eV. The modelling remains the same as before, except that the Stewart-Pyatt model [71] is used to model the effect of ionization potential depression on the bound-free energy. The DSFs are calculated for an incident beam energy of 9.0 keV and scattering angles of 30, 75, and 120, which are available using the Gbar experimental platform at the NIF.
[…]
The ray traced XRTS spectra are shown in Fig. 21 (b). A ray traced simulation of the elastic feature is also plotted to provide an indication of the shape and width of the SIF broadening. The FWHM of the feature is 91 eV, which is substantially larger than the $\sim 3$ eV broadening seen in the previous simulations.
The convergence of the inferred temperature for the three different ITCF ratios is plotted in Fig. 21 (c). As expected, the onset in convergence is much slower than for the previous experimental setup considered, requiring an integration limit $x\sim 400$–600 eV, depending on the ratio. The three different inferred temperatures are in good agreement with each other, but overestimate it to be 114–116 eV compared to the true value of 110 eV. That the temperature is reasonably close to the true value is perhaps surprising given the substantial variation in reflectivity across the spectral range; see Fig. 21 (a). Overall, this is a promising result that even very complicated instrument functions can be effectively accounted for within the ratios method. Additionally, convergence in the temperature is observed comfortably within the spectrometer’s spectral range.
@@ IV Summary and Discussion
Additionally, we have examined the applicability of the method to broad SIFs that are typical at laser facility experiments, such as the NIF. The integration range required to observe the temperature converge is large, but comfortably lies within the spectrometer spectral range. The specific spectrometer considered in those simulations was the MACS spectrometer, whose instrument function varies substantially within the considered spectral range, and so is not a simple convolution. Despite this, the ratios method is able to extract reasonable temperatures from the simulated spectra.
@@ Acknowledgments
This work was partially supported by the Center for Advanced Systems Understanding (CASUS), financed by Germany’s Federal Ministry of Education and Research (BMBF) and the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Union’s Just Transition Fund (JTF) within the project Röntgenlaser-Optimierung der Laserfusion (ROLF), contract number 5086999001, co-financed by the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2022 research and innovation programme (Grant agreement No. 101076233, “PREXTREME”). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Tobias Dornheim gratefully acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) via project DO 2670/1-1. A.K. and A.D.B. were partially supported by the U.S. Department of Energy Science Campaign 1. Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is co-authored by employees of NTESS. The employees, not NTESS, own the right, title and interest in and to the written work and are responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan. The work of M.P.B. was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344 and supported by Laboratory Directed Research and Development (LDRD) Grant No. 25-ERD-047. The authors gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy/Grete at NHR-Nord@Göttingen as part of the NHR infrastructure. The calculations for this research were conducted with computing resources under the project mvp00024.